OpenStudy (anonymous):

Calculate the flux of the vector field through the surface F=i +2j through a sphere of radius 3 at the origin

OpenStudy (experimentx):

perhaps we could apply divergence theorem ...

OpenStudy (anonymous):

hmmm that is next chapter

OpenStudy (experimentx):

you have the answer??

OpenStudy (anonymous):

nope not this time

OpenStudy (experimentx):

looks like we have to find tangential areal vector first.

OpenStudy (anonymous):

like shld i show u a previous example and how it was solved?

OpenStudy (experimentx):

yeah that would be helpful.

OpenStudy (anonymous):

ok so F=2i +3jthrough a disk of radius 5 in th eplane y=2oriented in the direction of increasing y

OpenStudy (anonymous):

Answer: area of disk is 25pi Flux = (21+3j)*25pij = 75pi

OpenStudy (experimentx):



OpenStudy (anonymous):

ummmm so?

OpenStudy (experimentx):

HA ... that was so nice and easy You multiplied Field (only y component) with area ...

OpenStudy (anonymous):

right but what do u do here

OpenStudy (anonymous):

when do i choose which component to use?

OpenStudy (experimentx):

 we know the area ... we don't know the component.

OpenStudy (anonymous):

right

OpenStudy (anonymous):

well in the example we knew the component and the area

OpenStudy (experimentx):

Formula was supposed to be surface integration of F.ds <--- and this ds is areal vector somehow I have a feeling that the answer will be zero.

OpenStudy (anonymous):

hmmm i may have the answer

OpenStudy (anonymous):

since F is a constant vector field the flux through a closed surface is 0 the flux that enters one side exits the other

OpenStudy (experimentx):

yeah ... but how to prove it ... without using divergence theorem.

OpenStudy (anonymous):

hmmm maybe i just need a better explanation or a diagram

OpenStudy (experimentx):

It would be something like this ... esp in case of sphere r^2 = x^2 + y^2 + z^2 Normal vector = del (r^2 - (x^2 + y^2 + z^2)) = -2xi - 2yj - 2zk the normal unit vectors are, 1/sqrt(3), 1/sqrt(3), 1/sqrt(3)

OpenStudy (experimentx):

sorry x/sqrt(3), y/sqrt(3), z/sqrt(3)

OpenStudy (experimentx):

Oo ... i forgot to put x's and y's down there x/rsqrt(3), y/rsqrt(3), z/rsqrt(3) , and since we already know r=3, we have x/3sqrt(3), y/3sqrt(3), z/3sqrt(3)

OpenStudy (anonymous):

ugghhh how did u get this equation?

OpenStudy (experimentx):

taking gradient of the equation of circle.

OpenStudy (anonymous):

ohhhh ok never did that b4

OpenStudy (experimentx):

really??? I thought you do these http://en.wikipedia.org/wiki/Del first. They are basics.

OpenStudy (anonymous):

hahahaha idk nemore

OpenStudy (experimentx):

okay still we can take standard unit normal vector as (x/3sqrt(3), y/3sqrt(3), z/3sqrt(3) and integrate it all over the surface of sphere.

OpenStudy (anonymous):

hey dont drive urself crazy. I needed the answer that i told u i guess iw ont prove it

OpenStudy (experimentx):

<(i +2j), (x/3sqrt(3), y/3sqrt(3), z/3sqrt(3))> = x/3sqrt(3) + 2y/3sqrt(3) well, if you say so then ..

OpenStudy (anonymous):

like ill just copy the answer like i am tooo tired to think like its 3 30 am here lol

OpenStudy (anonymous):

THANKS EXPERIMENT u r always a savior. like u never run off on me u always stay till i figure it out

OpenStudy (experimentx):

No ... it's rather I'm trying to learn things. anyway good night.

OpenStudy (anonymous):

Well thanks for ur time i really appreciate it